OpenStudy (anonymous):

Two equipotential surfaces surround a +2.90 x 10-8-C point charge. How far is the 220-V surface from the 85.0-V surface?

OpenStudy (dape):

The potential difference is \[ V=220-85=135. \] Now the potential V at distance d from a point charge Q is \[ V=\frac{Q}{4\pi\epsilon_0 d} \] If you solve this for d and plug in the potential difference and the charge you have your answer.

OpenStudy (anonymous):

im lost there but ill try any other info would of coarse be handy if you could

OpenStudy (anonymous):

can you tell me what e0 is

OpenStudy (dape):

It's the permittivity of vacuum, a physical constant. It's value is \[ \epsilon_0≈8.854188*10^{-12} \] It has units farads per meter.